A car and a plane travel 50 miles and 180 miles during the same time. The difference between their speeds is 143 miles per hour. Find the speed of the airplane.
Answer by Alan3354(69443)

d = rt
50 = rt
180 = (r+143)t
-----------
180 = rt + 143t
180 = 50 + 143t
t = 130/143 = 10/11 hours
r = d/t
-------
Car: 50/(10/11) = 55 mph
Plane: 180/(10/11) = 198 mph